OpenStudy (anonymous):

(x^(-2)-y^(-2))/(x^(-1)-y^(-1)) please?

OpenStudy (anonymous):

\[{x^{-2}-y^{-2} \over x^{-1}-y^{-1}}?\]

OpenStudy (amistre64):

\[\frac{x^{-2}-y^{-2}}{x^{-1}-y^{-1}}\] right?

OpenStudy (amistre64):

its just a matter of flipping all the ^- exponents to their reciprocals and re writeing it

OpenStudy (amistre64):

what part are you stuck at?

OpenStudy (anonymous):

\[x ^{-1}-y ^{-1}\]

OpenStudy (anonymous):

Just factor the numerator: \[={(x^{-1}-y^{-1})(x^{-1}+y^{-1}) \over x^{-1}-y^{-1}}=x^{-1}+y^{-1}\]

OpenStudy (anonymous):

so is it 1/x-y

OpenStudy (anonymous):

\[\frac{x^{-2}-y^{-2}}{x^{-1}-y^{-1}}=\frac{\frac{1}{x^2}-\frac{1}{y^2}}{\frac{1}{x}-\frac{1}{y}}\] Answar's ways method is the correct snappy way to do it but if you want to see why it is true, multiply numerator and denominator by \[x^2y^2\] to clear the fractions, or subtract and divide and you will get \[\frac{y-x}{xy}\]

OpenStudy (anonymous):

wow, thanks satelite73! how do i give a medal for your answer? you had the clearest solution of all! made me really understand the problem well. :) THANK YOU SO MUCH TO THE OTHERS AS WELL! <3

OpenStudy (anonymous):

By the way satellite, it's Anwar :) .. My way is simpler if you focus :P

OpenStudy (anonymous):

you're right Anwar! it is! i guess i just have to be more focused!